OpenStudy (anonymous):

I have to for what values of X is the expression undefined? x^2 - 24 / 3x^2 - 24x -60 I cant find any common factors, can someone help me work it out? By the calculator its -2 and 10

OpenStudy (anonymous):

In order for it to be undefined 3x^2-24x-60 has to be 0 at -2 and 10 that is the case so your calculator is right :)

OpenStudy (anonymous):

3x^2 - 24x -60 = 3(x^2 - 8x - 20) = 3(x - 10)(x + 2) this = 0 when x = 10 or x = -2 so expression is undefined for these values

OpenStudy (anonymous):

Thank you jimmyrep!! I worked it out right until the final step :)